2 kg of water at 23°C is placed in a refrigerator kept at 3°C. How much heat is lost by the water as it cools to 3°C? give your answer in joules. The answer should be a positive number.
Answer #1

Heat lost = mC(T1-T2)

m is mass, C is the specific heat capasity of water and is 4182 J/kg-k

Heat lost = 2 * 4182 * (23-3) = 167280 J